Circular No. 7682

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001do IN UGC 11459 M. Modjaz and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered Katzman Automatic Imaging Telescope CCD frames taken on Aug. 14.2 (mag ∼ 16.2) and 15.2 UT (mag ∼ 16.1). The new object is located at α = 19^{h}37^{m}23^{s}.02, δ = +40^{o}42′24′′.3 (equinox 2000.0), which is 5′′.2 east and 3′′.7 north of the nucleus of UGC 11459. A KAIT image taken on Aug. 10.2 showed nothing at this position (limiting mag ∼ 19.0).

SGR 1806–20 S. S. Eikenberry, M. Garske, D. Hu, M. A. Jackson, S. G. Patel, D. J.

Barry, M. Colonno, and J. R. Houck, Cornell University, report the discovery of possible infrared counterparts to the soft γ-ray repeater SGR 1806– 20: “Using archival Chandra observations, we find the x-ray counterpart to be located at α = 18^{h}08^{m}39^{s}.32, δ = −20^{o}24′39′′.8 (equinox 2000.0; uncertainty ± 0′′.5 in each coordinate) — in good agreement with the recently announced localization of Kaplan et al. (http://xxx.lanl.gov/abs/astroph/?0108195) based on longer Chandra observations. We took near-infra-red J, H, and K images of this field with the Cerro Tololo 4-m telescope (+ OSIRIS infrared camera) on July 6 UT under nonphotometric conditions, and photometrically calibrated them with more shallow observations of the field and an infrared standard star taken at the Hartung-Boothroyd Observatory 0.65-m telescope. Using eight USNO-A2.0 stars in the field for astrometric reference, we find an rms of 0′′.5 in the infrared astrometric solution, allowing us to locate the Chandra source to ∼ 1′′.0 accuracy, including systematic uncertainties in the Chandra astrometry. We find three possible counterparts within this error circle with K (2.2-μm) magnitudes of 16.2, 16.7, and ∼ 18, and with position end figures of 39^{s}.30, 40′′.5; 39^{s}.35, 38′′.5; and 39^{s}.32, 39′′.5, respectively. The two brightest objects are near the edges of the error circle, while the faintest is near the center. The brightest has color J −K = +2.6, making it likely to be a foreground object. Neither of the other objects is detectable in the J band, down to a 3σ limit of J > 21.0, and thus having color limits consistent with the massive star cluster near SGR 1806–20 (J − K ∼ +5.0). At the proposed distance and reddening towards SGR 1806–20 and the star cluster, the two fainter candidates have absolute K magnitudes of −2.3 and ∼ −1, respectively. Further observations to search for variability are needed to determine which, if either, of these candidates is the infrared counterpart to SGR 1806–20.”

2001 August 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green